# An open catalogue of digital world literature with likes

### v3

@@ Introduction
The use of text collections as a basis for linguistic inquiry has a long history that far predates modern computing. However, it was the computational turn of the late 1950s that established this practice as a fully-fledged academic discipline, later known as corpus linguistics. Dependent on technology capable of managing large quantities of machine-readable text [McEnery2003], early corpus linguistics often aimed to study the building blocks of language–e.g., grammar and lexis. With the increasing availability of computational power, other subfields soon broadened the practice’s aspirations. Sociolinguistics and discourse analysis are prominent illustrations of this methodological borrowing, applying corpus techniques to the study of social structures and norms. Similarly, the field of distant reading applies these methods to uncover patterns in literary history. In this context, the recent progression of Computational Social Science (css) can be seen as an advance in scale–utilising millions of books and social media data–and in tools–applying advanced machine learning and network analysis–rather than a fundamental change in the long-standing goal of understanding culture through texts.
@@
The vision of using extensive book corpora to represent culture at a large scale was significantly advanced by mass digitisation projects. The scanning effort undertaken by Google, beginning in 2002, culminated in the Google Books project and yielded, among other resources, the dataset for a foundational ‘quantitative analysis of culture’ [michel2011quantitative]. Shortly thereafter, the HathiTrust Digital Library [christenson2011hathitrust] was formed as a partnership between Google and major research libraries. Now containing over 18 million volumes, HathiTrust has become a go-to resource for making broad claims about culture through the computational analysis of books. Its utility is demonstrated by its application across diverse fields, including the history of science [murdock2017multi], literary history [underwood2019distant], gender studies [underwood2018transformation], and musicology [downie2020hathitrust].
@@
Despite its scale and utility, HathiTrust is not without significant limitations and biases, many of which stem from its origins. The collection itself is not a neutral representation of world literature; its composition privileges large research universities, while effectively excluding smaller, non-affiliated institutions and other forms of participation. Furthermore, as a collection built from scanned physical books, the quality of the underlying text can be inconsistent due to data integrity issues and errors from the Optical Character Recognition (ocr) process. The most significant challenge, however, relates to access. The ‘dark history’ [centivany2017dark] of HathiTrust’s formation was fraught with legal and political tensions surrounding copyright law. Consequently, a large portion of the collection containing in-copyright works remains inaccessible for direct reading. This forces researchers to operate within the controlled environment of their research centre’s secure ‘data capsule’ which not only presents a steeper learning curve but can also inhibit the broader, collaborative evolution of new computational tools and methods that thrive on open data. These combined issues of representational bias, data quality, and the constraints of a closed research environment highlight the challenges of using even the largest institutional corpora to study culture.
@@
These limitations have given rise to a new frontier for corpus-based research: the large-scale ‘shadow libraries’ that operate outside of formal institutions. Platforms like Library Genesis (LibGen), Sci-Hub, and Z-Library have emerged as a direct response to the access problem, aggregating tens of millions of books and articles in defiance of paywalls and copyright restrictions. These are not merely illicit collections; they are vast, user-driven archives whose very existence represents a form of crowd-sourced cultural curation, making them an essential new source of data for the computational social sciences and humanities [karaganis2018shadow]. The recent development of Large Language Models (llms) has opened a Pandora’s box in this regard, moving the use of these datasets from a niche practice to a central component of modern ai. Major companies have publicly acknowledged using such sources in academic papers [lu2024deepseek] and court cases [KadreyMeta2025, Anthropic2025], where their legal defence of ‘fair use’ has already been partially upheld.
@@ 1.1 Library Genesis, Z-Library and Anna’s archive
The shadow libraries that form the basis of our corpus have distinct but overlapping histories. The oldest and most foundational is Library Genesis (LibGen) [libgen2008www], which was started around 2008 to consolidate and share academic texts. Its origins are rooted in the clandestine ‘samizdat’ culture of the Soviet era and an ethos of providing free access to knowledge for academic communities facing economic hardship and institutional collapse. A pivotal moment in its history occurred in 2011, when LibGen absorbed the massive collection of the defunct shadow library Library.nu, transforming it from a primarily Russian-language archive into a global, multidisciplinary resource.
LibGen’s core operational principle is to be radically open, distributing not just its content but also its catalogue and source code to allow for a resilient, mirrored ecosystem. Appearing around the same time, Z-Library launched in 2009 and grew into one of the largest shadow libraries, with a collection that partially overlaps with LibGen’s but is separately administered and arguably less open [karaganis2018shadow]. Both platforms have faced significant legal challenges from publishers, culminating in Z-Library having several of its domains seized by the fbi in late 2022.
The most recent development in this ecosystem is Anna’s Archive [annarchive2022www], which appeared in 2022 and aims to provide a comprehensive, searchable index and mirror of other shadow libraries, including LibGen and Z-Library. For this study, we used LibGen’s own platform to access their data but relied on Anna’s Archive to source Z-Library’s content.
@@ 1.2 Discarding pdfs
N. Items
(in millions)
15.2 50.5 18.9 24.2
Herfindahl Index
0.24 0.74 0.32 0.57
@@ 2 Goodreads
Goodreads [goodreads2006www], a prominent social reading platform launched in 2006 and acquired by Amazon in 2013, boasts over 150 million members who rate, review, catalogue, and discuss books, creating a vast digital archive of contemporary reader responses and amateur criticism. Academics utilise this extensive dataset to analyse reading activity at scale, compare modern literary reception with historical patterns, and understand how works gain or lose popularity [walsh2021goodreads, antoniak2024afterlives, bourrier2020social, kousha2017goodreads, hu2025decides]. Nevertheless, these studies acknowledge limitations, including demographic biases within the user base (predominantly white, female, and us-centric) and the potential for review manipulation.
@@ 2.1 Data requirements and rationale
This Work-Edition scaffold partly relates to more advanced classification systems, for instance the Functional Requirements for Bibliographic Records (frbr) [funcReq1998], devised by the International Federation of Library Associations and Institutions (ifla). Their Work entity aligns closely with our eponymous category, both referring to a ‘distinct intellectual or artistic creation’. With a degree of interpretative flexibility–and in the specific context of books–our notion of edition corresponds roughly to the grouping of frbr’s Expression and Manifestation entities. This approach also conceptually parallels the method used by the Online Computer Library Center (oclc) to cluster bibliographic records into Work entities within WorldCat.
@@
Popularity is another common corpus selection criterion. Indeed, the very inclusion of a work in any corpus represents some kind of selection against obscurity, an escape from what Moretti termed ‘The Slaughterhouse of Literature’ [moretti2000slaughterhouse]. Different corpora embody distinct curation models: for instance, the HathiTrust dataset, composed of contributions from partner libraries, represents a decentralised yet expert-driven curation model, while shadow libraries are more crowd-sourced. Regardless of the base corpus, having further popularity metrics can prove useful to differentiate between popular and less popular books. Goodreads as a social network offers typical features, such as ratings and reviews. Integrating these features into our catalogue offers significant analytical potential, allowing researchers both to study popularity patterns and to select sub-corpora based on specific popularity criteria. This latter point is particularly relevant for computationally intensive ai research, where budget is a significant constraint. Popularity metrics allow a researcher to create an affordably-sized sub-corpus by sampling the most popular works, while still preserving the catalogue’s thematic and temporal representativeness.
@@
Before selecting Goodreads, we explored several alternative data sources. We believe it is relevant for future research to make our rationale for this choice explicit. The most significant alternative was Open Library [openlibrary2006www], a non-profit initiative of the Internet Archive, launched in 2006 with the ambitious goal of creating ‘one web page for every book ever published’.
@@ 2.2 The crawl of Goodreads
This iterative crawl continued until the author-led discovery of new items began to plateau at depth four and ceased at depth five, along with the complete exhaustion of the recommendation-led crawl. Both concomitant terminations suggest that our dataset is a comprehensive representation of Goodreads’ discoverable content, which resulted in our final dataset of 4,778,1244{,}778{,}124 works, 28,105,91328{,}105{,}913 editions, and 2,150,5222{,}150{,}522 authors (cf. Fig. 2). To the best of our knowledge, and based on our review of the state of the art [DBLP:conf/recsys/WanM18, DBLP:conf/acl/WanMNM19, JannesarGoodreads2020, BrightDataGoodreads2024, hu2025decides], this represents one of the largest publicly available crawls of Goodreads data published to date.
@@ 4 Legal and ethical considerations
At the time of writing this paper, legal frameworks in many countries are being adjusted to find a balance between copyright law and the ai use of copyrighted materials [sag2024globalization]. The legality of using shadow libraries is a central point of contention in recent lawsuits [KadreyMeta2025, Anthropic2025] and remains a contested, evolving issue. Jurisprudence continues to refine the relevant criteria for balancing these interests: the presence of social value from contributions to public research, the absence of market impact, and the fair use of a transformative application, to name only a few.
@@
MajinBook’s final output–a bibliographic index–is legally secure under the controlling us Supreme Court ruling in Feist Publications, Inc. v. Rural Telephone Service Co. [feist1991]. This case established that facts (such as the titles, authors, and publication dates in our catalogue) are not copyrightable. This same conclusion holds under the European Union’s Database Directive (96/9/EC) [databaseDirective1996].
MajinBook’s process–harvesting and computing data from shadow libraries–can be considered through the lens of several legal frameworks: In the eu, the Article 3 of the Directive on Copyright in the Digital Single Market (2019/790) [copyrightUEdirective2019], and in the us, the fair use doctrine incorporated into the Copyright Act of 1976 (17 usc § 107) [uscopyright1976] complemented by the recent Text and Data Mining exemption to the Digital Millennium Copyright Act (dmca) [uscfr_37CFR201_40]. The precise application of these frameworks is a complex and evolving debate, one that is beyond the scope of this paper and our expertise as non-legal scholars. However, from our reading, the core of the discussion–both legal and ethical–appears to hinge on the intent and context of the use. Our work situates in a public research institution, acting in good faith with no commercial interest and with the sole aim to foster public knowledge about language, literature and cultural representations.
@@
Third-party integration
To facilitate the diachronic analysis of these datasets, the books referenced in our primary English and secondary language corpora were integrated into the Gallicagram platform [^6] https://gallicagram-react.vercel.app/ by its developers. Gallicagram is an n-gram viewer designed to explore term frequencies across a wide collection of French and international datasets [azoulay2021gallicagram]. The tool also enables programmatic access to its underlying n-gram data via an api.

### v4

@@ Introduction
The use of text collections as a basis for linguistic inquiry has a long history that far predates modern computing. However, it was the computational turn of the late 1950s that established this practice as a fully-fledged academic discipline, later known as corpus linguistics. Dependent on technology capable of managing large quantities of machine-readable text [21], early corpus linguistics often aimed to study the building blocks of language–e.g., grammar and lexis. With the increasing availability of computational power, other subfields soon broadened the practice’s aspirations. Sociolinguistics and discourse analysis are prominent illustrations of this methodological borrowing, applying corpus techniques to the study of social structures and norms. Similarly, the field of distant reading applies these methods to uncover patterns in literary history. In this context, the recent progression of Computational Social Science (css) can be seen as an advance in scale–utilising millions of books and social media data–and in tools–applying advanced machine learning and network analysis–rather than a fundamental change in the long-standing goal of understanding culture through texts.
@@
The vision of using extensive book corpora to represent culture at a large scale was significantly advanced by mass digitisation projects. The scanning effort undertaken by Google, beginning in 2002, culminated in the Google Books project and yielded, among other resources, the dataset for a foundational ‘quantitative analysis of culture’ [22]. Shortly thereafter, the HathiTrust Digital Library [8] was formed as a partnership between Google and major research libraries. Now containing over 18 million volumes, HathiTrust has become a go-to resource for making broad claims about culture through the computational analysis of books. Its utility is demonstrated by its application across diverse fields, including the history of science [24], literary history [30], gender studies [29], and musicology [11].
@@
Despite its scale and utility, HathiTrust is not without significant limitations and biases, many of which stem from its origins. The collection itself is not a neutral representation of world literature; its composition privileges large research universities, while effectively excluding smaller, non-affiliated institutions and other forms of participation. Furthermore, as a collection built from scanned physical books, the quality of the underlying text can be inconsistent due to data integrity issues and errors from the Optical Character Recognition (ocr) process. The most significant challenge, however, relates to access. The ‘dark history’ [6] of HathiTrust’s formation was fraught with legal and political tensions surrounding copyright law. Consequently, a large portion of the collection containing in-copyright works remains inaccessible for direct reading. This forces researchers to operate within the controlled environment of their research centre’s secure ‘data capsule’ which not only presents a steeper learning curve but can also inhibit the broader, collaborative evolution of new computational tools and methods that thrive on open data. These combined issues of representational bias, data quality, and the constraints of a closed research environment highlight the challenges of using even the largest institutional corpora to study culture.
@@
These limitations have given rise to a new frontier for corpus-based research: the large-scale ‘shadow libraries’ that operate outside of formal institutions. Platforms like Library Genesis (LibGen), Sci-Hub, and Z-Library have emerged as a direct response to the access problem, aggregating tens of millions of books and articles in defiance of paywalls and copyright restrictions. These are not merely illicit collections; they are vast, user-driven archives whose very existence represents a form of crowd-sourced cultural curation, making them an essential new source of data for the computational social sciences and humanities [17]. The recent development of Large Language Models (llms) has opened a Pandora’s box in this regard, moving the use of these datasets from a niche practice to a central component of modern ai. Major companies have publicly acknowledged using such sources in academic papers [20] and court cases [26, 2], where their legal defence of ‘fair use’ has already been partially upheld.
@@ 1.1 Library Genesis, Z-Library and Anna’s archive
The shadow libraries that form the basis of our corpus have distinct but overlapping histories. The oldest and most foundational is Library Genesis (LibGen) [19], which was started around 2008 to consolidate and share academic texts. Its origins are rooted in the clandestine ‘samizdat’ culture of the Soviet era and an ethos of providing free access to knowledge for academic communities facing economic hardship and institutional collapse. A pivotal moment in its history occurred in 2011, when LibGen absorbed the massive collection of the defunct shadow library Library.nu, transforming it from a primarily Russian-language archive into a global, multidisciplinary resource.
LibGen’s core operational principle is to be radically open, distributing not just its content but also its catalogue and source code to allow for a resilient, mirrored ecosystem. Appearing around the same time, Z-Library launched in 2009 and grew into one of the largest shadow libraries, with a collection that partially overlaps with LibGen’s but is separately administered and arguably less open [17]. Both platforms have faced significant legal challenges from publishers, culminating in Z-Library having several of its domains seized by the fbi in late 2022.
The most recent development in this ecosystem is Anna’s Archive [3], which appeared in 2022 and aims to provide a comprehensive, searchable index and mirror of other shadow libraries, including LibGen and Z-Library. For this study, we used LibGen’s own platform to access their data but relied on Anna’s Archive to source Z-Library’s content.
@@ 1.2 Discarding pdfs
N. Items (in millions) 15.2 50.5 18.9 24.2
Herfindahl Index (normalized) 0.24 0.74 0.32 0.57
@@ 2 Goodreads
Goodreads [7], a prominent social reading platform launched in 2006 and acquired by Amazon in 2013, boasts over 150 million members who rate, review, catalogue, and discuss books, creating a vast digital archive of contemporary reader responses and amateur criticism. Academics utilise this extensive dataset to analyse reading activity at scale, compare modern literary reception with historical patterns, and understand how works gain or lose popularity [31, 4, 5, 18, 16]. Nevertheless, these studies acknowledge limitations, including demographic biases within the user base (predominantly white, female, and us-centric) and the potential for review manipulation.
@@ 2.1 Data requirements and rationale
This Work-Edition scaffold partly relates to more advanced classification systems, for instance the Functional Requirements for Bibliographic Records (frbr) [25], devised by the International Federation of Library Associations and Institutions (ifla). Their Work entity aligns closely with our eponymous category, both referring to a ‘distinct intellectual or artistic creation’. With a degree of interpretative flexibility–and in the specific context of books–our notion of edition corresponds roughly to the grouping of frbr’s Expression and Manifestation entities. This approach also conceptually parallels the method used by the Online Computer Library Center (oclc) to cluster bibliographic records into Work entities within WorldCat.
@@
Popularity is another common corpus selection criterion. Indeed, the very inclusion of a work in any corpus represents some kind of selection against obscurity, an escape from what Moretti termed ‘The Slaughterhouse of Literature’ [23]. Different corpora embody distinct curation models: for instance, the HathiTrust dataset, composed of contributions from partner libraries, represents a decentralised yet expert-driven curation model, while shadow libraries are more crowd-sourced. Regardless of the base corpus, having further popularity metrics can prove useful to differentiate between popular and less popular books. Goodreads as a social network offers typical features, such as ratings and reviews. Integrating these features into our catalogue offers significant analytical potential, allowing researchers both to study popularity patterns and to select sub-corpora based on specific popularity criteria. This latter point is particularly relevant for computationally intensive ai research, where budget is a significant constraint. Popularity metrics allow a researcher to create an affordably-sized sub-corpus by sampling the most popular works, while still preserving the catalogue’s thematic and temporal representativeness.
@@
Before selecting Goodreads, we explored several alternative data sources. We believe it is relevant for future research to make our rationale for this choice explicit. The most significant alternative was Open Library [28], a non-profit initiative of the Internet Archive, launched in 2006 with the ambitious goal of creating ‘one web page for every book ever published’.
@@ 2.2 The crawl of Goodreads
This iterative crawl continued until the author-led discovery of new items began to plateau at depth four and ceased at depth five, along with the complete exhaustion of the recommendation-led crawl. Both concomitant terminations suggest that our dataset is a comprehensive representation of Goodreads’ discoverable content, which resulted in our final dataset of 4,778,1244{,}778{,}124 works, 28,105,91328{,}105{,}913 editions, and 2,150,5222{,}150{,}522 authors (cf. Fig. 2). To the best of our knowledge, and based on our review of the state of the art [32, 33, 14, 15, 16], this represents one of the largest publicly available crawls of Goodreads data published to date.
@@ 4 Legal and ethical considerations
At the time of writing this paper, legal frameworks in many countries are being adjusted to find a balance between copyright law and the ai use of copyrighted materials [27]. The legality of using shadow libraries is a central point of contention in recent lawsuits [26, 2] and remains a contested, evolving issue. Jurisprudence continues to refine the relevant criteria for balancing these interests: the presence of social value from contributions to public research, the absence of market impact, and the fair use of a transformative application, to name only a few.
@@
MajinBook’s final output–a bibliographic index–is legally secure under the controlling us Supreme Court ruling in Feist Publications, Inc. v. Rural Telephone Service Co. [13]. This case established that facts (such as the titles, authors, and publication dates in our catalogue) are not copyrightable. This same conclusion holds under the European Union’s Database Directive (96/9/EC) [12].
MajinBook’s process–harvesting and computing data from shadow libraries–can be considered through the lens of several legal frameworks: In the eu, the Article 3 of the Directive on Copyright in the Digital Single Market (2019/790) [10], and in the us, the fair use doctrine incorporated into the Copyright Act of 1976 (17 usc § 107) [9] complemented by the recent Text and Data Mining exemption to the Digital Millennium Copyright Act (dmca) [1]. The precise application of these frameworks is a complex and evolving debate, one that is beyond the scope of this paper and our expertise as non-legal scholars. However, from our reading, the core of the discussion–both legal and ethical–appears to hinge on the intent and context of the use. Our work situates in a public research institution, acting in good faith with no commercial interest and with the sole aim to foster public knowledge about language, literature and cultural representations.